OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−5, 5) and a directrix of y = -1.

OpenStudy (anonymous):

(-5,5)

OpenStudy (anonymous):

f(x) = −one twelfth (x − 5)2 + 2 f(x) = one twelfth (x − 5)2 + 2 f(x) = −one twelfth (x + 5)2 + 2 f(x) = one twelfth (x + 5)2 + 2

OpenStudy (anonymous):

all those diamond things are supposed to be ( - ), i don't know why they show up like that

OpenStudy (anonymous):

thought it was the third one but it was wrong too

hero (hero):

Given two points, the focus (-5,5) and the directrix (x, -1), plug both points into this formula: \((x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\) Then simplify

OpenStudy (anonymous):

\[(-5 - x)^2 + (5 - 1)^2 = (-5 - x)^2 + (5 - 1)^2 \] What did I do wrong here? I come up with zero

hero (hero):

You're only supposed to plug the values for \((x_1, y_1)\) and \((x_2, y_2)\)

hero (hero):

\((x_1, y_1)\) = (-5, 5) \((x_2, y_2)\) = (x, -1)

OpenStudy (anonymous):

Ok so (x - x1)^2 + (y - y1)^2 = (x - x2)^2 + (y - y2)^2 (x - 5)^2 + (y - 5)^2 = (x - x)^2 + (y - 1)^2 Is it like this \[(x - 5)^2 + (y - 5)^2 = (x - x)^2 + (y - 1)^2 \] or this \[(x + 5)^2 + (y - 5)^2 = (x - x)^2 + (y + 1)^2\] I feel like I'm plugging these in wrong. :/

hero (hero):

\((x + 5)^2 + (y - 5)^2 = (x - x)^2 + (y + 1)^2\) is correct. Try to understand why.

OpenStudy (anonymous):

Because two ( - ) make a ( + )

hero (hero):

Okay, now what can we do next?

OpenStudy (anonymous):

simplify

hero (hero):

Simplify what?

OpenStudy (anonymous):

Umm the expression? or it doesn't need simplifying?

hero (hero):

It has to be simplified. I was just asking about your simplification plan? How are you going to approach simplifying this?

hero (hero):

@CrazyTurtle483, you still here?

OpenStudy (anonymous):

Yes sorry my dad wanted me to go get something

OpenStudy (anonymous):

Ok honestly i really tried to simplify this but i dont understand where to start \[(x + 5)^2 + (y - 5)^2 = (x - x)^2 + (y + 1)^2 \]

OpenStudy (anonymous):

+ x or do a +5?

hero (hero):

Start with \((x - x)^2\) Obviously x - x = 0

OpenStudy (anonymous):

makes sense

hero (hero):

and 0^2 = 0

hero (hero):

Next, subtract \((y - 5)^2\) from both sides

OpenStudy (anonymous):

(x + 5)^2 + (y - 5)^2 = (y + 1)^2 - (y - 5)^2 - (y - 5)^2 (x + 5)^2 = 12y − 24?

OpenStudy (anonymous):

\[(x + 5)^2 = 12y - 24\] I don't know why there are diamonds

hero (hero):

Show me how you got from \((y + 1)^2 - (y - 5)^2\) to 12y - 24

OpenStudy (anonymous):

I didn't know what (y+1)^2 - (y-5)^2 was so i distributed and combined like terms

hero (hero):

What do you mean? Explain your algebraic steps.

OpenStudy (anonymous):

since youre asking about 12y -24, is that wrong? i dont know what (y+1)^2 - (y-5)^2 equals.. distribute y^2 + 2y + 1 + − y^2 + 10y + −25 then you combine the like terms (2y + 10y) + (y2 - y2) + (1 + -25) 12y + -24

hero (hero):

Actually, instead of expanding what you do is use the difference of squares formula

OpenStudy (anonymous):

\[y^2 + 2y + 1 + −y^2 + 10y + −25 \]

OpenStudy (anonymous):

i havent learned that yet

hero (hero):

Are you saying you don't know how to factor \(y^2 - 4\) ?

OpenStudy (anonymous):

(y + 2) (y - 2) ?

hero (hero):

Exactly I know that you know the difference of squares

OpenStudy (anonymous):

didn't know factoring was also called difference of squares haha

OpenStudy (anonymous):

ok so \[(x+5)^2 = (y+2)(y-2) ? \]

hero (hero):

Difference of Squares is \(a^2 - b^2 = (a + b)(a - b)\)

hero (hero):

\(y^2 - 4 = y^2 - 2^2 = (y + 2)(y - 2)\)

hero (hero):

\((y + 1)^2 - (y - 5)^2 = (y + 1 + y - 5)(y + 1 - (y - 5))\)

hero (hero):

and \((y + 1 + y - 5))(y + 1 - (y - 5)) = (2y - 4)(6) = 12y - 24\)

hero (hero):

Either way, 12y - 24 is correct

hero (hero):

Now all you have to do is just continue solving for y

OpenStudy (anonymous):

ok so would it be 6(2y - 4) = 12y - 24

OpenStudy (anonymous):

then what because don't you distribute next?

hero (hero):

Isolate y

hero (hero):

What you have left is \((x + 5)^2 = 12y - 24\)

OpenStudy (anonymous):

ok but then what, isn't y isolated? or 12y - 24 y = 2

hero (hero):

You don't know how to isolate y from there?

hero (hero):

Add 24 to both sides. Afterwards, divide both sides by 12

OpenStudy (anonymous):

(x+5)^2 = 12y - 24 +24 +24 ok so youre left with 12y one one side but what is (x+5)^2 + 24?

OpenStudy (anonymous):

i feel like i should know this better

OpenStudy (anonymous):

\[\frac{ 1 }{ 12 } (x+5)^{2} + 2 ? \]

OpenStudy (anonymous):

you there @Hero

hero (hero):

Looks right

OpenStudy (anonymous):

Yay it was!

OpenStudy (anonymous):

thank you so much man, you are one of the most patient people ive ever met

hero (hero):

We don't stop until it's done and right.

OpenStudy (anonymous):

Most people would probably have given up on me haha, thanks again.

hero (hero):

No, it wasn't that bad.